OpenStudy (anonymous):

The length of a rectangle is three more times than twice its width. Determine the dimensions that will give a total area of 27 m ^2.

OpenStudy (studygurl14):

"The length of a rectangle is three more times than twice its width." L = 2W + 3 "Area of 27" LW = 27 Take those two equations and solve for the variables.

OpenStudy (wcrmelissa2001):

(just clarifying for you) L is the length of the rectangle W is the width of the rectangle So your first equation is L=2W+3 And second is LW=27 So substitute in your value of L from your first equation iNTO the second equation: (2W+3)*W = 27 Solve from there